jowhite:

What is the force required to accelerate 1 kilogram of mass at 1 meter per second per second. 1 pound 1 newton 1 kilometer 1 gram

new2luv:

f=ma

new2luv:

so what's the mass

new2luv:

so the mass is 1 and the acceleration would be 1 so do f=1*1 and you get 1 newton

new2luv:

@jowhite make sense

jowhite:

Thx

new2luv:

yw

jhonyy9:

@new2luv wrote:
f=ma
F = m×a F - force m - mass a - acceleration

new2luv:

@jhonyy9 wrote:
@new2luv wrote:
f=ma
F = m×a F - force m - mass a - acceleration
yes